Status and Runplan, Sunday, August 26

From HallCWiki
Jump to: navigation, search
  • Status
    • Tune beam in Hall C - optics work underway at start of swing shift 8/26
    • MQE3M03 correct polarity?
    • IHA3H07A - no wire traces - need to contact EES
  • Run plan
    • Sunday SWING, Monday OWL to DAY
      • Beam optics tuning
      • After initial setup - SHMS steering plan will be executed (please call RC if not already on site). When MCC asks, ramp up magnets (one at a time) to settings for -6.053 GeV.
        • Note: It is possible that the steering test may not be completed before end of swing (or before Jay leaves). If so, over OWL shift, proceed with the beam and detector checkout using HMS only. The SHMS magnets may not be turned on until the steering test is complete.
      • Beam checkout
      • For the following, see detailed run plan here: [1]
        • Detector and trigger checkout. Call experts before start: Malace, Pooser, Berdnikov, Ambrose for detector checkout, Malace, Poose, Huber, Trotta for trigger checkout
        • Calibration runs